OpenStudy (alex_mattucci):

Please Help Me!!! Need help with some polynomial problems. Marvel Man movie is being released to a limited number of movie theaters. The number of showtimes for the movie is twice the number of movie theaters. The number of seats available for each showtime is three times the number of showtimes. Let x represent the number of theaters showing the movie. A. Determine the three factors that are to be multiplied to determine the total number of seats available for the movie. B. When simplified what is the degree and classification of this expression? C. Interpret each pairing of factors as a single unit in the context of the problem.

OpenStudy (alex_mattucci):

Please help @SolomonZelman

Directrix (directrix):

First part courtesy of @ranga Number of theaters showing the movie = x "number of showtimes for the movie is twice the number of movie theaters" Number of showtimes for the movie = 2 * x "number of seats available for each showtime is three times the number of showtimes" Number of seats available for each showtime = 3 * 2 * x = 6x The three factors that need to be multiplied to determine the TOTAL number of seats are: 6x, 2x and x. http://openstudy.com/updates/52f1903fe4b0d693ffe3a5f8

Directrix (directrix):

B. When simplified what is the degree and classification of this expression? Multiply these: 6x, 2x and x. What is the degree of the resulting monomial? @Alex_Mattucci

OpenStudy (alex_mattucci):

First Degree Trinomial? @Directrix

OpenStudy (alex_mattucci):

Ok thank you so much @Directrix With your help I Figured it out!!!!

Directrix (directrix):

B. When simplified what is the degree and classification of this expression? If these 3 factors are multiplied: 6x, 2x and x. the product is 12*x^3 . That is a single term so it is a monomial. The degree is 3 because the variable x is raised to the third power. @Alex_Mattucci

OpenStudy (alex_mattucci):

Oh yeah, I see that, Thanks. I am going to change that real quick!

OpenStudy (alex_mattucci):

Just wondering, could I bother you for help with one more @Directrix ?

Directrix (directrix):

Yes. I am not sure about this 3rd part, though. I have not worked on it. C. Interpret each pairing of factors as a single unit in the context of the problem.